### v2

# SciLitLLM: How to Adapt LLMs for Scientific Literature Understanding

Abstract.
Scientific literature understanding is crucial for extracting targeted information and garnering insights, thereby significantly advancing scientific discovery. Despite the remarkable success of Large Language Models (LLMs), they face challenges in scientific literature understanding, primarily due to (1) a lack of scientific knowledge and (2) unfamiliarity with specialized scientific tasks.
To develop an LLM specialized in scientific literature understanding, we propose a hybrid strategy that integrates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse scientific domain knowledge and enhance instruction-following capabilities for domain-specific tasks. In this process, we identify two key challenges: (1) constructing high-quality CPT corpora, and (2) generating diverse SFT instructions. We address these challenges through a meticulous pipeline, including PDF text extraction, parsing content error correction, quality filtering, and synthetic instruction creation. Applying this strategy, we present a suite of LLMs: SciLitLLM, specialized in scientific literature understanding. These models demonstrate promising performance on scientific literature understanding benchmarks. Specifically, the 7B model shows an average performance improvement of 3.6% on SciAssess and 10.1% on SciRIFF compared to leading LLMs with fewer than 15B parameters. Additionally, the 72B model, trained using QLoRA, achieves state-of-the-art performance among widely adopted open-source models.
Our contributions are threefold: (1) We present an effective framework that integrates CPT and SFT to adapt LLMs to scientific literature understanding, which can also be easily adapted to other domains. (2) We propose an LLM-based synthesis method to generate diverse and high-quality scientific instructions, resulting in a new instruction set – SciLitIns – for supervised fine-tuning in less-represented scientific domains. (3) SciLitLLM achieves promising performance improvements on scientific literature understanding benchmarks. We release the data processing codes [^1] https://github.com/dptech-corp/Uni-SMART and model weights [^2] https://huggingface.co/Uni-SMART/SciLitLLM.
Large Language Model, Pre-training, Supervised Fine-tuning, Scientific Literature Understanding
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing Methodologiesccs: Artificial Intelligenceccs: Natural Language Processingccs: Natural Language Generation
1. Introduction
Figure 1. An example of scientific literature understanding in SciRIFF involves extracting accurate entities from a biomedicine paper. SciLitLLM-7B demonstrates sufficient scientific knowledge and instruction-following ability to accurately identify and extract these entities.
Scientific literature understanding involves the systematic evaluation and interpretation of scientific texts and publications, to identify trends, extract targeted information, and garner insights (AI4Science and Quantum, 2023; Zheng et al., 2023), significantly contributing to scientific discovery. Concurrently, Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Yang et al., 2024; Touvron et al., 2023) have achieved remarkable success in natural language processing, prompting the development of domain-specific LLMs across various fields (Cui et al., 2024; Wu et al., 2023; Clusmann et al., 2023). However, recent studies (Wadden et al., 2024; Cai et al., 2024; Singh et al., 2023) indicate that LLMs face challenges when specializing in scientific literature understanding, particularly in context understanding and question answering. Take Figure 1 as an example, where the LLM is asked to understand the content of a biomedical research paper and then extract the targeted information. LLMs’ potential might be hindered by two major barriers: (1) a lack of scientific knowledge, which results in errors such as the missing important entities in Llama-3-7B (AI@Meta, 2024), and (2) unfamiliarity with scientific tasks, leading to the inability of Galactica-6.7B (Taylor et al., 2022) to follow task instructions accurately.
To make LLMs specialized in science-relevant tasks, existing studies mostly adopt two strategies, as illustrated in Figure 2: (1) Fine-tuning with scientific instructions (Wadden et al., 2024; Zhang et al., 2024b; Singhal et al., 2022; Huang et al., 2023). A general-purpose LLM is fine-tuned with collected domain-specific instructions to adapt it to science-relevant tasks. However, instruction fine-tuning alone is insufficient to imbue the models with comprehensive scientific knowledge. (2) Pre-training on scientific corpora (Beltagy et al., 2019; Zeng et al., 2022; Taylor et al., 2022). This approach involves training models on vast scientific corpora. While this method equips LLMs with domain knowledge, the lack of instruction-tuning confines them to solving relevant tasks. Moreover, it is hampered by substantial computational costs and data requirements (Ling et al., 2023; Yang et al., 2023). To address these obstacles while balancing efficiency, we propose a hybrid strategy that incorporates continual pre-training (CPT) and supervised fine-tuning (SFT), to simultaneously infuse domain knowledge and enhance domain-specific instruction-following capabilities.
Figure 2. Comparison of strategies to adapt LLMs to scientific tasks. Previous approaches typically either fine-tune a general LLM with scientific instructions or pre-train an LLM on extensive scientific corpora. We propose a combined method of both CPT and SFT.
However, as illustrated in Figure 3, developing a scientific literature understanding model using this CPT and SFT pipeline presents two critical requirements:
  • High-quality CPT Corpora. Scientific corpora, predominantly in PDF format such as textbooks and research papers, are not directly digestible for LLM training. Converting these documents to text using tools like PyPDF2333https://pypdf2.readthedocs.io often introduces formatting and syntax errors, degrading corpus quality. Worse still, scientific documents often contain segments that contribute little information (e.g., references), necessitating quality control to filter them out. See the first row in Figure 3 for a comparison of high- and low-quality CPT texts.
  • Diverse Scientific Instructions. Effective instruction following for scientific literature understanding requires a large, high-quality, and diverse set of task-related instructions. However, to the best of our knowledge, there is a scarcity of well-designed instruction datasets for scientific literature understanding, and hiring human annotators to curate such a dataset from scratch is prohibitively expensive (Erdmann et al., 2019; Qiu et al., 2023). See the second row in Figure 3 for an illustration of high- and low-quality instructions.
To address these challenges, we devise an effective pipeline to construct high-quality domain corpora for CPT and diverse scientific instructions for SFT, as illustrated in Figure 4:
  • In the CPT stage for domain knowledge injection, we start with an extensive in-house corpus consisting of 73k textbooks and 625k academic papers in the scientific field, all in PDF format. Initially, we leverage PyPDF2, a widely used open-source PDF parsing tool, to extract raw texts from these documents. We then employ a moderate yet powerful model, Llama3-7B-Instruct (AI@Meta, 2024), to correct the format and spelling errors introduced by PDF parsing (cf. Section 3.1.1). Subsequently, we train a small text quality classifier to score the corpus and filter out texts of low educational value [^4] Phi models (Gunasekar et al., 2023; Li et al., 2023a; Abdin et al., 2024) propose to determine the quality of a pre-training text by its educational value for a student whose goal is to learn basic domain concepts. in the scientific field (cf. Section 3.1.2). These two simple yet effective textual refinement and quality control measures ensure the high quality of our CPT corpus, culminating 12.7 billion tokens for CPT via the Qwen2 (Yang et al., 2024) tokenizer.
  • In the SFT stage for domain instruction fine-tuning, to overcome the scarcity of domain-specific instructions and the high cost of human annotations, we propose a novel instruction synthesis method (cf. Section 3.2.1). It enables us to generate diverse instructions to better equip the model for domain-specific tasks. Moreover, we sequentially apply instruction duplication based on Levenshtein distance and an LLM-based filtering method to ensure the quality of synthetic instructions (cf. Section 3.2.2).
Figure 3. Examples of high and low-quality CPT text and SFT instructions. Scores are labeled by the CPT and SFT quality filter (cf. Section 3.1.2 &  3.2.2), ranging from 0 to 5. Higher scores indicate better quality.
Having established such high-quality datasets, we apply the CPT-SFT integration strategy on a general-purpose LLM – Qwen2 (Yang et al., 2024) and obtain SciLitLLM in two scales: a 7B (full parameter training) and a 4-bit quantized 72B (QLoRA (Dettmers et al., 2023) parameter-efficient training) model. Evaluations on benchmarks of scientific literature understanding demonstrate the effectiveness of our strategy. We observe promising performance enhancements, with an average improvement of 3.6% on SciAssess (Cai et al., 2024) and 10.1% on SciRIFF (Wadden et al., 2024), compared to the leading LLMs under 15B parameters. Notably, SciLitLLM-7B even outperforms Llama3 and Qwen2 with 70B parameters on SciRIFF. Additionally, SciLitLLM-72B, with QLoRA training, achieves leading results on both benchmarks, surpassing other open-source LLMs. Further ablation studies demonstrate the effectiveness of each module in our pipeline.
In summary, our contributions are threefold:
  • We devise an effective and comprehensive pipeline to adapt general LLMs to a specific domain – scientific literature understanding. It combines continual pre-training (CPT) and supervised fine-tuning (SFT), to enhance scientific knowledge base and instruction-following capabilities for specialized domain tasks.
  • We propose a novel domain instruction synthesis method to curate instructions for scientific literature understanding, resulting in a new dataset – SciLitIns .
  • SciLitLLM, trained through the proposed pipeline, outperforms leading open-source LLMs on scientific literature understanding. It has been deployed in real-world application scenarios and has demonstrated promising performance. Code and data will be released after administrative procedure.
2. Related Works
In this section, we provide a review of literature related to continual pre-training, supervised fine-tuning in the scientific domain, and LLMs for scientific literature understanding.
2.1. Knowledge Injection via Continual Pre-training
Pre-training a language model is usually conducted on a large corpus of textual data to learn the statistical properties of language (Radford et al., 2019; Brown et al., 2020). Formally, given a sequence of textual tokens $x=(x_{1},x_{2},...,x_{T})$, an LLM parameterized by $\theta$ performs autoregressive language modeling (ALM) by predicting the next token in the sequence given the previous tokens:
(1) $\mathcal{L}_{\text{ALM }}(\theta)=\sum_{t=1}^{T}\log P_{\theta}\left(x_{t}\mid x_{<t}\right).$
To further inject domain knowledge into a general LLM after pre-training, researchers engage in continual pre-training (CPT) on high-quality domain-specific corpora (Jin et al., 2022; Sun et al., 2020), sometimes combined with general corpora. This process enhances the model’s fundamental understanding abilities in specific downstream domains while mitigating catastrophic forgetting of general knowledge (Wu et al., 2022; Ke et al., 2023; Mehta et al., 2023). See the comprehensive study (Gupta et al., 2023) for different warm-up strategies for CPT. Additionally, the CPT corpora can be augmented by transforming them into an instruction-response format (Cheng et al., 2023, 2024). Furthermore, the scaling law (Hoffmann et al., 2022) of domain-specific CPT (Que et al., 2024) is explored to determine the optimal mix of data. However, these studies primarily focus on training dynamics and data recipes, leaving the pre-processing for scientific data, especially raw PDF files, largely unexplored. Exhibiting such steps is essential for generating high-quality domain corpora and effectively injecting domain knowledge, representing a significant challenge for practitioners.
2.2. Domain Adaptation via Supervised Fine-tuning
Supervised fine-tuning (SFT) modifies a pre-trained language model to follow specific instructions or perform designated tasks by fine-tuning it on a targeted, task-specific dataset (Raffel et al., 2020; Wei et al., 2021). Let $\mathcal{D}_{\text{fine}}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ represent the fine-tuning dataset, where $x_{i}$ is an input sequence (instruction) and $y_{i}$ is the corresponding output sequence. The fine-tuning objective can be expressed as:
(2) $\mathcal{L}_{\text{fine-tune }}(\phi)=\sum_{i=1}^{M}\log P_{\phi}\left(y_{i}\mid x_{i}\right).$
During fine-tuning, the model adjusts its parameters $\phi$ to better fit the task-specific data, typically involving parameter-efficient (Hu et al., 2021; Dettmers et al., 2023) or full parameter training (Raffel et al., 2020; Wei et al., 2022a). Applying SFT to a general LLM for specific domain adaptation has demonstrated effectiveness in various fields: in medicine (Clusmann et al., 2023), corpora of medical literature and clinical notes are used; in law (Cui et al., 2024), legal documents and case law are compiled; and in finance (Wu et al., 2023), financial reports and market data are utilized. In the scientific domain, several studies have specialized LLMs for scientific tasks, often necessitating the construction of a substantial domain-specific dataset with SFT. For example, SciGLM (Zhang et al., 2024a) leverages existing LLMs to generate step-by-step reasoning for unlabelled scientific instructions. ChemLLM (Zhang et al., 2024b), a more specified LLM in the chemistry field, collects structured chemical data from a vast selection of online databases and transforms this structured data into a question-answering format for SFT. SciRIFF (Wadden et al., 2024) converts existing literature understanding datasets into natural language input-output pairs suitable for instruction-tuning using pre-defined templates. However, benchmark studies (Feng et al., 2024; Cai et al., 2024) indicate that SFT alone may not provide adequate scientific knowledge to excel in relevant tasks. This suggests the need for a more comprehensive approach that combines domain knowledge infusion with instruction-following enhancements.
Figure 4. The pipeline of SciLitLLM consists of two key stages: continual pre-training (CPT) for scientific knowledge injection and supervised fine-tuning (SFT) for scientific instruction following. Specifically, the CPT stage involves several modules: PDF parsing, format & grammar correction (cf. Section 3.1.1), and quality filtering (cf. Section  3.1.2) modules. These modules ensure the model is equipped with high-quality scientific domain knowledge. The SFT stage includes LLM-based instruction generation (cf. Section 3.2.1) and instruction quality control (cf. Section 3.2.2) measures. These steps are designed to fine-tune the model’s ability to follow scientific instructions accurately and effectively.
2.3. LLMs for Scientific Literature Understanding
In the scientific domain, existing strategies for developing specialized LLMs mostly fall into two categories: (1) Supervised fine-tuning with scientific instructions. This approach requires a large, high-quality, and diverse set of instructions to cultivate problem-solving abilities for scientific tasks. Representative works (e.g., SciGLM (Zhang et al., 2024a), ChemLLM (Zhang et al., 2024b), and SciRIFF (Wadden et al., 2024)) have been detailed in Section 2.2. (2) Pre-training with scientific corpora. This approach involves pre-training on a large corpus of scientific texts to improve performance on downstream scientific tasks. Early attempts, such as SciBert (Beltagy et al., 2019) and KV-PLV (Zeng et al., 2022), are based on BERT (Devlin et al., 2019) and pre-trained on a large corpus of scientific text for downstream scientific task enhancement. More recently, Galactica (Taylor et al., 2022) is pre-trained on a vast corpus of scientific literature, including research papers, scientific articles, and other relevant scientific texts. Despite these advances, two major limitations hinder these models from excelling in scientific literature understanding: (1) lack of scientific knowledge, and (2) inability to follow task instructions. To address these challenges, we propose a combined pipeline of CPT and SFT to devise a specialized LLM for scientific literature understanding. It injects domain-specific knowledge through CPT while enhancing task-specific instruction-following abilities through SFT, leading to a more capable LLM for scientific literature understanding.
3. Method
In this section, we discuss the details of our proposed pipeline (cf. Figure 4): continual pre-training for scientific knowledge injection (cf. Section 3.1) and supervised fine-tuning for scientific tasks enhancement (cf. Section 3.2).
3.1. CPT for Scientific Knowledge Injection
What are high-quality pre-training corpora? Researchers (Gunasekar et al., 2023; Li et al., 2023a; Abdin et al., 2024) suggest that language models benefit from corpora that possess the same qualities as an exemplary textbook for human learners: clarity, self-containment, instructiveness, and balance. These characteristics ensure that the material is not only comprehensible but also informative and comprehensive, providing a solid foundation for knowledge acquisition. Over the past decades, the efforts of scientists and educators have resulted in a wealth of high-quality scientific textbooks and research papers, which serve as invaluable resources for learning and teaching. Recognizing this, we have curated a substantial collection of over 73,000 textbooks and 625,000 research papers within the scientific domain, ensuring all documents are copyright-compliant. To inject their rich scientific knowledge into a general LLM, we perform continual pre-training (CPT) on these high-quality textbooks and papers. This process equips the model with a robust scientific knowledge base, thereby paving the way for developing a specialized LLM tailored for scientific literature understanding.
However, we still face two practical obstacles when dealing with those textbooks and research papers: (1) Formatting and syntax errors. Most textbooks and research paper documents are in PDF format, which is not directly digestible by LLMs. Consequently, we need to transform them into plain text. Converting these documents using tools like PyPDF2 often introduces formatting and syntax errors, which degrade the quality of the corpus. (2) Corpus quality control. Despite their overall high quality, textbooks and research papers also contain segments with little useful information, such as references and garbled text introduced during the PDF parsing process. Given the large scale of the pre-training corpora, an effective and computation-efficient quality control measure is essential.
To tackle these obstacles, we devised the following modules of format & grammar correction and CPT quality filter:
3.1.1. Format & Grammar Correction
As illustrated in Appendix A, a parsed text from a PDF document often contains many formatting and syntax errors. To address this issue, we prompt a moderate yet powerful language model, Llama3-7B-Instruct (AI@Meta, 2024), to correct these errors introduced during the PDF parsing process. Utilizing the vLLM (Kwon et al., 2023) backend, Llama3-7B-Instruct can process approximately 2.52 million tokens per Nvidia A100 GPU hour. The process takes over 5,000 A100 GPU hours to handle all 73,000 textbooks and 625,000 research papers. Example texts – both before and after processing – along with the prompt template are provided in Appendix A to demonstrate the improvements made through this correction process.
3.1.2. CPT Quality Filter
During CPT, maintaining the quality of the training corpus is crucial for effective knowledge injection. Given the extremely large scale of pre-training corpora, assessing quality through human annotation is not feasible (Erdmann et al., 2019; Qiu et al., 2023). Consequently, leading LLMs (e.g., Phi (Gunasekar et al., 2023), Llama (Touvron et al., 2023), and Qwen (Yang et al., 2024)) employ model-based quality filters. The typical process involves using larger LLMs to score the quality of a subset of texts, which then serve as labels for training small classifiers (e.g., random forest (Gunasekar et al., 2023) and Bert (AI@Meta, 2024)) to annotate the entire training corpus. Inspired by this approach, we design a resource-efficient method based on a lightweight text quality classifier.
Following prior studies (Gunasekar et al., 2023; AI@Meta, 2024; Yang et al., 2024; Anton et al., 2024), we first annotate a random subset of 50k CPT texts using a powerful model – Llama3-70B-Instruct (AI@Meta, 2024). We adapt the quality assessment prompt from fineweb-edu-classifier (Anton et al., 2024), a widely-used quality classifier for web data, to evaluate the educational value (Gunasekar et al., 2023) of the scientific knowledge in each sampled text, assigning scores ranging from 0 (lowest quality) to 5 (highest quality). After annotation, we perform supervised transfer learning on the fineweb-edu-classifier (Anton et al., 2024) checkpoint – a Bert-based (Devlin et al., 2019) quality classifier in the web domain. This process results in a scientific text quality classifier tailored for scientific corpus assessment. See Appendix B for more details about classifier training and hyperparameters.
We then utilize this classifier to assess the quality of the entire CPT dataset (See Figure 3 for concrete samples). Each sample is evaluated and assigned with a continuous real number as the quality score. To enhance the overall quality of training data, we then exclude the lowest-scoring 25% from the dataset.
By leveraging the CPT quality classifier, we can efficiently filter out low-quality texts and ensure that only high-quality, informative content is retained. This step is crucial for enhancing the scientific knowledge base of our LLM, thereby improving its performance in scientific literature understanding.
3.1.3. CPT Training Settings
We perform CPT on Qwen2-Base (Yang et al., 2024) for one epoch, encompassing 23.7 billion tokens (cf. Table 1), with a sequence length of 2,048 tokens. To maintain the model’s general knowledge, we also include a similar scale of general corpus tokens from Redpajama (Computer, 2023). To stabilize the learning procedure, we gradually decrease the learning rate from $1\times 10^{-5}$ to 0 for SciLitLLM-7B, and from $5\times 10^{-6}$ to 0 for SciLitLLM-72B (QLoRA), with a cosine scheduler. To address overfitting, we apply a weight decay of $0.1$ and gradients were clipped at a maximum value of $1.0$. The CPT training took approximately 3 days on 32 Nvidia A100 GPUs for SciLitLLM-7B-Base (full parameters) and about 10 days for SciLitLLM-72B-Base (QLoRA).
Stage Data source Domain #Doc/# Ins # Tokens
CPT In-house Textbooks Science 73k 10B
In-house Journals Science 625k 2.7B
Redpajama (Computer, 2023) General - 11B
SFT SciLitIns Science 110k 86M
SciRIFF (Wadden et al., 2024) Science 70k 40M
Infinity-Instruct [^5] https://huggingface.co/datasets/BAAI/Infinity-Instruct General 3M 1.7B
Table 1. Data statistics of continual pre-training and supervised fine-tuning. #Doc/#Ins denotes the number of documents of CPT corpora and the number of instructions for SFT, respectively. Underlined datasets are curated by us.
3.2. SFT for Scientific Instruction Following
After performing CPT on an extensive scientific corpus to incorporate domain knowledge, we subsequently conduct SFT on domain-specific instructions to enhance the model’s ability to understand scientific literature. We identify two major challenges in SFT for scientific instruction following:
  • Existing instruction-tuning datasets in the scientific domain (Feng et al., 2024; Fang et al., 2023; Li et al., 2024) primarily focus on fields such as physics, chemistry, and biology. Manually collecting instruction-tuning data for other less-represented vertical domains (e.g., alloy, biomedicine, and material) is both time-consuming and costly (Erdmann et al., 2019; Qiu et al., 2023).
  • Few instruction-tuning datasets adequately reflect the scenario of scientific literature understanding, which typically involves a segment of scientific literature accompanied by a question that requires deriving an answer from the text.
To address these challenges, we draw inspiration from leading models (e.g., Nemotron-4 (Adler et al., 2024), Phi (Gunasekar et al., 2023), and Qwen (Yang et al., 2024)), which leverage existing LLMs to construct synthetic instruction sets. We propose a novel instruction synthesis method to curate instructions specifically for scientific literature understanding.
3.2.1. Instruction Synthesis of Less-represented Domains
Unlike typical question-answer pairs, an instruction for a scientific literature understanding task comprises three components: (1) a segment of scientific literature, (2) a question pertaining to the context, and (3) the corresponding answer (Wadden et al., 2024). Simply prompting an LLM to generate a scientific context along with an associated question-answer pair – without variations in the instructions or parameters – often yields similar or repeated contents. This phenomenon arises because language models tend to adhere to the most probable or common paths dictated by their memory base and priors, thereby lacking the creativity to explore diverse generation (Gunasekar et al., 2023). Consequently, we are motivated to devise a strategy to encourage the language model to produce more creative and diverse instructions for scientific literature understanding while simultaneously ensuring the quality and coherence of the generated contexts.
We design a simple yet effective three-step pipeline to generate diverse and high-quality instructions for scientific contexts and corresponding question-answer pairs, consisting of the following:
  1. (1)
    Probability table of domain keywords. For a target scientific domain (e.g., alloy, biomedicine, and material), we collect dozens of high-impact research papers via Google Scholar [^6] https://scholar.google.com/ and count the frequency of each word appearing in these papers. Subsequently, we remove spaces and meaningless articles such as “a,” “an,” “the,” etc. After normalization, a probability table of domain keywords, representing the word-level distribution of domain literature, is obtained.
  2. (2)
    Scientific task descriptions. Since LLMs are expected to handle various types of scientific tasks, an instruction set with task diversity is essential. Therefore, we compile a list of task descriptions by including representative tasks from existing scientific NLP datasets (Wadden et al., 2024; Cai et al., 2024; Feng et al., 2024), covering as many scenarios as possible that an LLM may encounter in real applications.
  3. (3)
    Instruction Generation. Given a word probability table and the task list for a specific scientific domain, we sample 20 keywords and a task description each time. Subsequently, GPT-4o (OpenAI, 2023) is prompted to generate a scientific context containing the sampled keywords and a question-answer pair according to the provided task description.
The detailed generation process and example prompts are presented in Appendix C.1. Utilizing this pipeline, we obtain over 100k synthetic instructions for scientific literature understanding, covering less-represented scientific domains and various types of specialized tasks.
3.2.2. Instruction Quality Control
Figure 5. The quality of SciLitInsis evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability (the higher the better). Instructions with an average score of less than 4 are filtered out.
To ensure the diversity and quality of generated instructions, effective measures for quality control are essential. Specifically, we incorporate heuristic deduplication and LLM-based filtering.
  1. (1)
    Heuristic deduplication. Despite the measures taken during the generation process to prevent high homogeneity in the instructions, the generated data points may still contain similar questions or identical answers. To eliminate such redundancy, we implement a simple yet effective deduplication process using the Levenshtein distance to calculate the similarity score between instructions. Based on this score, $5\%$ to $10\%$ of similar data are removed for each question type. Detailed processing steps are provided in Appendix C.2.
  2. (2)
    LLM-based filtering. Inspired by recent efforts (Eldan and Li, 2023; Chiang and Lee, 2023; Zhang et al., 2024c) to measure the quality of generated content using LLMs, we leverage Llama-3-70B-Instruct (AI@Meta, 2024) to assess the quality of generated instructions for scientific literature understanding. Specifically, the quality is evaluated from five aspects: clarity, complexity, correctness, usefulness, and adaptability, assigning each instruction a score from 0 (lowest quality) to 5 (highest quality). Instructions with an average score of less than 4 are filtered out. We show the quality statistics of synthetic instructions in Figure 5. The detailed recipe for instruction quality evaluation with concrete examples is included in Appendix C.3.
Through instruction synthesis and quality control pipeline, we obtain SciLitIns, consisting of approximately 110,000 high-quality and diverse instructions for scientific literature understanding. With these instructions, models’ problem-solving abilities in this specialized field could be enhanced.
3.2.3. SFT Training Settings
Our SFT training dataset consists of three parts: SciLitIns, SciRIFF (Wadden et al., 2024) and Infinity-Instruct [^7] https://huggingface.co/datasets/BAAI/Infinity-Instruct, as shown in Table 1. Infinity-Instruct is a collection of more than twenty open-source instructions datasets, covering various general domains. SciRIFF and SciLitIns contain specialized instructions for scientific literature understanding. We use full parameter training for SciLitLLM-7B-Base and QLoRA (Dettmers et al., 2023) parameter-efficient training for SciLitLLM-72B-Base. For both models, we train for one epoch on Infinity-Instruct to cultivate their general instruction-following abilities, then for five epochs on SciLitIns and SciRIFF for scientific literature understanding enhancement. The training is conducted with a sequence length of 4,096, a maximum learning rate of $5\times 10^{-6}$, and a cosine scheduler. The SFT training takes approximately 32 hours on 32 A100 GPUs for the 7B and 100 hours for the 72B model, resulting in SciLitLLM-7B-Instruct and SciLitLLM-72B-Instruct.
Models MMLU-s CMMLU-s Xiezhi-en-s Xiezhi-ch-s
# Parameters ¡ 15B
ChatGLM-6B 46.99 46.42 58.33 63.50
Mistral-7B 59.84 36.43 64.97 53.34
Qwen1.5-7B 57.09 73.74 65.87 71.60
Qwen2-7B 66.62 86.04 71.53 74.36
Llama2-7B 39.87 28.33 42.40 36.40
Llama3-8B 61.52 42.16 66.29 63.46
Llama2-13B 49.06 32.29 58.30 42.65
Qwen1.5-14B 66.67 80.19 68.60 73.94
SciLitLLM-7B 70.85 91.84 73.42 78.24
# Parameters ¿ 50B
Mixtral-8x7B 67.58 44.88 69.54 65.81
Qwen2-57B-A14B 73.79 89.65 70.35 73.73
Llama2-70B 65.03 43.94 66.75 65.98
Llama3-70B 76.43 66.41 71.36 73.28
Qwen1.5-72B 74.89 86.87 71.18 74.47
Qwen2-72B 80.86 92.31 72.41 75.03
Qwen1.5-110B 77.06 90.72 73.45 73.14
SciLitLLM-72B (QLoRA) 82.31 93.08 74.23 76.93
Table 2. Performance comparison of base models. Bold indicates the highest performance for LLMs under 15B parameters or above 50B parameters. SciLitLLM achieves leading performance on all four scientific knowledge benchmarks.
Dataset Domain/ Task # Parameter ~7B # Parameter ~70B API
SciTulu-7B Mistral-7B Llama3-8B Qwen2-7B SciLitLLM-7B Llama3-70B Qwen2-72B SciLitLLM-72B (QLoRA) GPT3.5 GPT4o
SciAssess FundSci 32.3 48.3 58.5 70.3 74.8 70.9 77.1 78.4 62.2 76.7
AlloyMat 23.9 28.0 32.9 32.8 35.6 44.9 42.7 49.1 32.0 52.1
Biomed 67.8 76.0 77.4 80.8 79.6 79.6 81.0 79.6 78.0 82.3
DrugDisc 25.4 30.2 32.0 31.7 33.2 41.5 35.5 41.8 31.0 43.4
OrgMat 16.7 20.6 24.5 28.3 38.9 41.5 52.7 48.6 24.4 62.7
[…]
SciRIFF BioASQ 37.5 43.9 44.7 40.7 51.0 46.3 43.6 50.7 47.3 46.7
BioR 55.7 48.2 45.3 44.3 74.0 59.9 59.1 63.0 53.9 61.0
DiscMT 61.5 44.6 58.7 59.9 77.4 71.9 73.5 73.5 67.9 78.3
EI 11.6 17.1 14.7 14.4 22.3 22.0 24.1 23.1 19.2 24.7
MC 34.6 47.0 49.5 51.6 68.0 59.9 57.3 70.5 47.8 58.7
MuP 72.1 93.4 90.7 96.6 76.8 96.4 97.8 77.9 76.8 86.9
Qasper 54.2/38.6 58.6/39.4 58.2/41.9 56.8/34.5 58.4/56.9 25.0/19.4 63.3/47.2 60.5/54.1 54.7/39.8 67.8/50.5
SciERC 35.6 30.2 19.9 27.5 39.9 35.2 34.1 46.9 28.6 42.2
SciFact 66.0/49.2 68.5/51.3 64.6/51.7 65.2/44.3 68.5/59.7 85.1/67.3 82.3/65.9 75.2/60.6 69.7/53.3 84.3/68.7
[…]
Table 3. Model performances on scientific literature understanding benchmarks: SciAssess and SciRIFF. SciLitLLM-7B and SciLitLLM-72B achieve leading performance compared with models of similar scales. The best-performing models in the 7B and 70B scales are highlighted in bold. Results for SciTulu-7B, GPT-3.5, and GPT-4o on SciRIFF are taken from its original papers, while all other results are generated by our experiments.
4. Experiments
In this section, we perform experiments to answer the following research questions:
  • RQ1: How does SciLitLLM perform on scientific literature understanding tasks?
  • RQ2: Can CPT with domain-specific corpora aid in scientific knowledge injection?
  • RQ3: Can SFT with synthetic instructions improve performance on scientific literature understanding tasks?
4.1. Experimental Setup
4.1.1. Benchmarks
To evaluate the performance of LLMs regarding scientific knowledge base and specialized task-solving abilities, our benchmarks include:
  • CPT benchmarks. We evaluate the base models on three widely adopted benchmarks: MMLU (Hendrycks et al., 2021), CMMLU (Li et al., 2023b), and Xiezhi (Gu et al., 2024). Specifically, we select the STEM subsets from these benchmarks to assess their scientific knowledge, which serves as the foundation for scientific literature understanding.
  • SFT benchmarks. We evaluate the instruct models on scientific literature understanding benchmarks: SciRIFF (Wadden et al., 2024) and SciAssess (Cai et al., 2024). Brief descriptions of them are provided in Appendix D.
4.1.2. Baselines
We test the following baselines:
  • CPT baselines: We compare SciLitLLM-base against leading open-source base models: ChatGLM (Zeng et al., 2024), Llama3 (AI@Meta, 2024), Llama2 (Touvron et al., 2023), Qwen2 (Yang et al., 2024), Qwen1.5 (Team, 2024), Mistral-7B (Jiang et al., 2023) and Mixtral-8x7B (Jiang et al., 2024).
  • SFT baselines: We benchmark leading instruction LLMs including GPT-4o (OpenAI, 2023), GPT-3.5 (Brown et al., 2020), Llama3 (AI@Meta, 2024) and Qwen2 (Yang et al., 2024). We also report the performance of SciTulu-7B (Wadden et al., 2024), which is a fine-tuned Llama2-7B (Touvron et al., 2023) on SciRIFF.
4.2. Performance Overview (RQ1)
4.2.1. Base model performance
The performance comparison of base models is shown in Table 2. SciLitLLM-base consistently outperforms other general base models across four scientific benchmarks. Specifically, compared with LLMs of less than 15 billion parameters, SciLitLLM-7B-Base shows an average accuracy improvement of 3.9% over Qwen2-7B. For LLMs with more than 50 billion parameters, SciLitLLM-72B-Base, with QLoRA training, outperforms all other LLMs (without quantization) as large as 110 billion parameters. The results demonstrate the effectiveness of CPT on high-quality scientific corpora, paving the way to a specialized LLM for scientific literature understanding.
4.2.2. Instruct model performance
As shown in Table 3, SciLitLLM-7B-Instruct achieves the highest performance in 4 out of 5 domains on SciAssess, outperforming the second-best model by 3.6%. Notably, on SciRIFF, it surpasses baseline models by a substantial margin of 10.1%. Additionally, SciLitLLM-72B, trained using QLoRA, shows a 1.7% and 0.9% performance improvement over Qwen2-72B on SciAssess and SciRIFF, respectively.
Detailed model performance on SciAssess is presented in Table 7, where SciLitLLM-7B and SciLitLLM-72B both lead in 12 and 13 out of 29 sub-tasks. Specifically, SciLitLLM-7B excels in tasks such as table extraction and molecule generation, likely benefiting from the comprehensive task coverage in our synthetic instruction dataset, SciLitIns. On SciRIFF, SciLitLLM-7B/SciLitLLM-72B ranks first in 8/6 out of 11 evaluations [^8] In SciRIFF, the Qasper and SciFact tasks have two different evaluation methods and thus two results..
4.3. Ablation Study (RQ2 & RQ3)
We conducted ablation experiments on three key components in our pipeline: the CPT stage, the SFT data recipe, and the instruction quality filter, to demonstrate their effectiveness. It is important to note that all ablation experiments were performed on SciLitLLM-7B due to budget constraints.
4.3.1. Scientific knowledge injection via CPT (RQ2)
Model SciAssess SciRIFF
Qwen2-7B-Instruct 48.8 48.7
Qwen2-7B-Base + SFT 48.1 51.6
Qwen2-7B-Base + CPT + SFT 52.4 59.4
Table 4. Ablation study of the CPT stage. The results demonstrate the effectiveness of the CPT stage in improving performance on scientific literature understanding.
We investigate the contribution of the CPT stage for SciLitLLM. We compare the three variants: (1) Qwen2-7B-Instruct: official instruct-model checkpoint; (2) Qwen2-7B-base + SFT: applying our SFT stage directly to Qwen2-7B-base without CPT; (2) Qwen2-7B-base + CPT + SFT: SciLitLLM-7B-Instruct.
As shown in Table 4, applying SFT alone to the Qwen2-7B-Base model does not lead to clear performance gains on SciAssess (-0.7%) and yields only a modest improvement on SciRIFF (+2.9%). In contrast, incorporating both CPT and SFT results in substantial performance enhancements: a 3.6% improvement on SciAssess and a 10.7% gain on SciRIFF. These results demonstrate that the CPT is crucial for effectively injecting scientific knowledge and significantly enhancing LLM performance on scientific literature understanding tasks.
4.3.2. Influence of SFT Data Recipes (RQ3)
SFT Dataset SciAssess SciRIFF
Infinity-Instruct 44.5 44.7
Infinity-Instruct + SciRIFF 42.2 53.9
Infinity-Instruct + SciRIFF + SciLitIns 52.4 59.4
Table 5. Ablation study of SFT data recipes.
We explore the influence of each ingredient in SFT data recipes. We incrementally add three datasets to the SFT training set: Infinity-Instruct, SciRIFF, and SciLitIns. As shown in Table 5, using only the Infinity-Instruct  results in the lowest performance on both SciAssess (44.5%) and SciRIFF (44.7%). This indicates that fine-tuning LLMs on general instructions alone is insufficient for scientific literature understanding, likely because Infinity-Instruct lacks specialized contents.
Adding SciRIFF to Infinity-Instruct improves performance on SciRIFF significantly but decreases performance on SciAssess. This discrepancy may be due to the disjoint coverage of scientific domains between SciRIFF and SciAssess. Finally, including SciLitIns along with Infinity-Instruct and SciRIFF boosts performance on both benchmarks, with SciAssess at 52.4% and SciRIFF at 59.4%. This demonstrates that including SciLitIns that covers less-represented scientific domains and tasks is beneficial for enhancing model performance in scientific literature understanding.
4.3.3. Influence of Instruction Quality Filter
Dataset SciAssess SciRIFF
SciLitIns w/o filtering 51.1 56.2
SciLitIns w/ filtering 52.4 59.4
Table 6. Ablation study of SFT instruction quality filtering.
We conduct an ablation study to assess the impact of quality filter for synthetic instructions by varying whether the dataset SciLitIns  was filtered. As discussed in Section 3.2.2, this filter removes low-quality instructions evaluated from five key aspects. Table 6 shows that applying the filter significantly improved the performance of SciLitLLM-7B on SciAssess (+1.3%) and SciRIFF (+3.2%). This demonstrates that SFT quality filtering process effectively selects high-value educational instructions, thereby boosting the performance of SciLitLLMon scientific literature understanding.
5. Limitations
Despite the promising results achieved by SciLitLLM, there are several limitations that should be acknowledged:
  • Insufficient data volume. Compared with existing pre-training datasets (AI@Meta, 2024; Yang et al., 2024; Taylor et al., 2022), the amount of data used for CPT is not satisfying. Future work should consider incorporating a larger scientific corpus, potentially including scientific blogs or purely synthetic data, to enhance the model’s scientific knowledge base and overall performance.
  • Lack of reasoning enhancement. The current pipeline does not explore advanced reasoning techniques such as Chain-of-Thought (Wei et al., 2022b) or Tree-of-Thought (Yao et al., 2023) in the data construction or model inference stages. Investigating these methods could potentially improve the model’s inference capabilities and overall performance.
  • Lack of preference alignment. Due to a limited financial budget, the model lacks Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022). RLHF has shown significant improvements in aligning models with human preferences and ensuring more reliable outputs. Implementing RLHF in future iterations could further enhance the model’s reliability.
Addressing these limitations in future research will be crucial developing a more robust and capable LLM specialized in scientific literature understanding.
6. Conclusion and Future Works
In this paper, we introduce SciLitLLM, a specialized model for scientific literature understanding. It is initialized with a general base model – Qwen2 (Yang et al., 2024), and trained through a sequential pipeline of continual pre-training (CPT) and supervised fine-tuning (SFT). For effective scientific knowledge injection during CPT, we propose model-based format and grammar correction method, along with text quality filtering measures. To ensure high-quality and diverse instructions during SFT, we devise instruction synthesis and quality control approaches. Our experiments on widely-used benchmarks demonstrate the effectiveness of this pipeline in adapting a general model to the field of scientific literature understanding. Specifically, SciLitLLM-7B achieves a 3.6% improvement on the SciAssess (Cai et al., 2024) and a 10.1% improvement on the SciRIFF (Wadden et al., 2024) compared to leading models with fewer than 10 billion parameters. SciLitLLM-72B, trained with QLoRA, also surpasses baseline open-source LLMs. We note that this pipeline could be easily adapted to other specialized domains, particularly those lacking adequate open-source corpora and high-quality instruction sets.
Our future work will focus on expanding the diversity and quality of the training data, as well as exploring more efficient methods for domain-specific knowledge injection and high-quality instruction generation. Moreover, we plan to expand our pipeline to include the RLHF (Ouyang et al., 2022) stage for better human preference alignment and enhanced safety.
Appendix A Format & Grammar Correction Examples
Raw text parsed by PyPDF2 Highly p e n e t r a t i n g radiation, such as $\gamma$-rays or fast electorns, deposits ener gy throughout the solid t a r g e t material. Gas production occurs w i t h i n the solid phase and must d i f f u s e to the surface to be observed. The a p p a r e n t yield of H 2 can depend on the radiolysis pro c e d u r e or the particle size if some of the gas remains in the solid. Experiments have shown that the apparent y i e l d of H 2 can vary by a f a c t o r of 3 in the r a d i o l y s i s of polyethylene spheres of 7 to 2100 cm2/g (about 9 to 0.03 mm) [12]. The e f f e c t s of gas trapping and diffusion are not understood in the c o n t e x t of waste storage. Extremely h i g h dose rates in the p r o c e s s i n g of certain materials may lead to bubble formation, which could a l t e r product quality. The y i e l d of H 2 in the r a d i o l y s i s of polymers w i t h $\gamma$-rays is well known for several types of p o l y m e r s [2].
Format & Grammar Correction Prompt I have extracted the following raw text from a PDF, but the extraction process has introduced many formatting issues such as unnecessary line breaks, extra spaces, and other artifacts that disrupt the text flow. Could you please help me correct these formatting issues and provide a clean, readable version of the text? Respond with the Corrected Version only. Raw Text: {RawText} Start your response with ”Here is the corrected version of the text:”.
Text after Format & Grammar Correction Highly penetrating radiation, such as $\gamma$-rays or fast electrons, deposits energy throughout the solid target material. Gas production occurs within the solid phase and must diffuse to the surface to be observed. The apparent yield of H2 can depend on the radiolysis procedure or the particle size if some of the gas remains in the solid. Experiments have shown that the apparent yield of H2 can vary by a factor of 3 in the radiolysis of polyethylene spheres of 7 to 2100 cm2/g (about 9 to 0.03 mm) [12]. The effects of gas trapping and diffusion are not understood in the context of waste storage. Extremely high dose rates in the processing of certain materials may lead to bubble formation, which could alter product quality. The yield of H2 in the radiolysis of polymers with $\gamma$-rays is well known for several types of polymers [2].
Appendix B CPT Quality Filter
We randomly select 50k samples from our 700k CPT data. These selected samples are then scored using the Llama3-70B model. The prompt utilized for this scoring process is as follows:
Prompt for CPT Data Quality Labelling Below is an extract from a textbook. Evaluate whether the text has a high educational value and could be useful in an educational setting for teaching from primary school to grade school levels using the additive 5-point scoring system described below. Points are accumulated based on the satisfaction of each criterion:
- Add 1 point if the extract provides some basic information relevant to educational topics, even if it includes some irrelevant or non-academic content like advertisements and promotional material. - Add another point if the extract addresses certain elements pertinent to education but does not align closely with educational standards. It might mix educational content with non-educational material, offering a superficial overview of potentially useful topics, or presenting information in a disorganized manner and incoherent writing style. - Award a third point if the extract is appropriate for educational use and introduces key concepts relevant to school curricula. It is coherent though it may not be comprehensive or could include some extraneous information. It may resemble an introductory section of a textbook or a basic tutorial that is suitable for learning but has notable limitations like treating concepts that are too complex for grade school students. - Grant a fourth point if the extract is highly relevant and beneficial for educational purposes for a level not higher than grade school, exhibiting a clear and consistent writing style. It could be similar to a chapter from a textbook or a tutorial, offering substantial educational content, including exercises and solutions, with minimal irrelevant information, and the concepts aren’t too advanced for grade school students. The content is coherent, focused, and valuable for structured learning. - Bestow a fifth point if the extract is outstanding in its educational value, and perfectly suited for teaching either at primary school or grade school. It follows detailed reasoning, the writing style is easy to follow, and offers profound and thorough insights into the subject matter, devoid of any non-educational or complex content.
After examining the extract: - Briefly justify your total score, up to 100 words. - Conclude with the score using the format: ”Educational score: ¡total points¿”
We train a Scientific Texts Quality Classifier on these labeled data samples. The classifier is a 109M BERT (Devlin et al., 2019) classifier, fine-tuned from the checkpoint of fineweb-edu-classifier (Anton et al., 2024). The model is trained for 20 epochs with a learning rate of 0.001 and a batch size of 1024. Ninety percent of the 50K samples are used as the training set, and the rest 10% are used as the validation set. The training process costs approximately 50 minutes on 4 A100 GPUs. We select the checkpoint from the epoch that yields the highest validation micro F1 score as our final checkpoint.
During Inference, we set batch size to 2048, and beam number to 1. The inference process costs 90 minutes on 4 A100 GPUs. We utilize the generated to filter out 25% data with the lowest quality. The distribution of the scores is demonstrated in Figure 6. The filtered-out 25% data are marked gray, while the remaining 75% CPT data are marked orange.
Figure 6. Score distribution of the CPT Data
Appendix C SFT Details
C.1. Instruction Generation Pipeline
In SciLitIns, we focus on generating instructions for three less-represented domains (materials science, medicine, and drug discovery) and five question types:
  • Table Extraction: Table Extraction tasks evaluate a model’s proficiency in extracting, summarizing, and structuring data from an article into a table format.
  • Entity Extraction: Entity Extraction tasks are designed to evaluate a model’s ability to extract specific information, such as entities or relationships, from the text.
  • Molecule Translation: Molecule Translation tasks evaluate a model’s ability to translate molecules between different SMILES formats.
  • Molecule Extraction: Molecule Extraction tasks ask a model to extract an appropriate molecule from a scientific paragraph that contains multiple molecules.
  • Multiple Choice and True-or-False: Multiple Choice and True-or-False questions assess a model’s ability to select the correct answer from a set of options, testing its knowledge and reasoning on both simple and complex scenarios.
For each of the three scientific domains, we collect a set of high-impact research papers and construct a word frequency table. To generate a question in a given domain, we sample 20 keywords from the corresponding word table and insert them into the prompt for that question. To ensure fair representations of less frequent keywords, we use random sampling with a temperature setting of 3. We release our code, prompt templates, and word frequency tables. Below is an example of generating a table extraction question:
Prompt for Generating a Table Extraction Question I need synthetic training data for training a machine learning model that extracts tables from text correctly. The data should be formatted in JSON, with each entry containing ”text” and ”answer” attributes. You should generate a paragraph that includes the keywords: {{keywords}}. The ”text” part must contain enough information for the table to be extracted! In ”text” part, You must you include a table description in latex format. Special notice for the table content: You should generate a table that has complicated numbers and characters, include non-standard characters, and have a variety of values. Make sure the value you generated do not follow simple patterns, for example, never include deplicate values or values with constant interval in columns. Your answer should contain as much details as possible. You should only generate one JSON. The value for the two attributes should be two string. Use {{ and }} to warp your output. Pay attention to the escape characters in the latex format. Remember to put a comma at the end of the first string. Never use a json block to wrap your output. Here is the format for your output: { ”text”: ”Your paragraph here, remember to include a table in latex format”, ”answer”: ”Your answer table here” } Now start your answer:
C.2. Instruction Deduplication
The generated synthetic data may contain similar questions or identical answers. To eliminate redundancy, we implement a fuzzy deduplication process using the Levenshtein distance to calculate the similarity score between question-answer pairs. Specifically, for two pairs $(q_{1},a_{1})$ and $(q_{2},a_{2})$, their textual similarity is defined as $(1-\text{lev}(q_{1},q_{2}))(1-\text{lev}(a_{1},a_{2}))$, where $\text{lev}(\cdot,\cdot)$ denotes the Levenshtein distance. Due to significant differences between texts from different question types, we compute similarity matrices separately for each type. We then use a disjoint-set data structure to merge highly similar data points. We use this process to remove approximately 5% to 10% of duplicated data for each question type.
C.3. Quality Assessment of Generated SFT Instructions
In section 3.2.2, we sample 10k instruction pairs from SciLitIns and evaluate them by Llama-3-70B using the below prompt.
[…]
Below is an example of SciLitIns, which will be sent to Llama-3-70B for evaluation.
[…]
Instruction
Extract the table in the paragraph.
Response
Material,Initial Enthalpy (kJ/mol),Final Enthalpy (kJ/mol),Fracture Phase Material A,25.4,47.8,Brittle Material B,22.3,45.0,Ductile Material C,28.9,50.2,Semi-brittle
Below is the evaluation by Llama-3-70B on the example from  SciLitIns. Each criterion is scored out of 5 points, making the total possible points 25. The total score is then averaged, resulting in an overall score out of 5.
Evaluation by Llama-3-70B Explanation: The instruction is clear and specific, asking the model to extract a table from a given paragraph. The paragraph itself is well-written, with proper grammar and syntax, making it easy to understand. The table is also well-structured and provides valuable information about enthalpy changes during fracture phases. The task requires a good understanding of the text and the ability to identify the table. However, the complexity of the task is relatively low, as it only involves extracting a table, and the context is not particularly nuanced or ambiguous. The task is useful for advancing knowledge in material science, and the table could be adapted to different contexts or use cases. ==================== ”Clarity”: 5, ”Complexity”: 2, ”Correctness”: 5, ”Usefulness”: 4, ”Adaptability”: 4, ”Total”: 20
Appendix D Benchmark Details
To the best of our knowledge, there are two commonly-adopted datasets for scientific literature understanding:
  • SciRIFF (Wadden et al., 2024) evaluates essential scientific literature understanding capabilities, including information extraction, summarization, question answering, claim verification, and classification. Data points in SciRIFF are notable for their long input contexts and complicated structured outputs. The Qasper and SciFact tasks have two different evaluation methods and thus two results. We note that SciRIFF contains a separate training set used in the SFT stage in our study.
  • SciAssess (Cai et al., 2024) features an end-to-end benchmark of understanding PDF content. It includes 29 tasks from five scientific domains: fundamental science, alloy materials, biomedicine, drug discovery, and organic materials. SciAssess is used exclusively for testing in our evaluation.
Overall, SciRIFF provides basic benchmarks for comprehending short scientific segments and various instructions, while SciAssess presents more challenging tasks involving longer contexts from raw PDFs.
Appendix E Detailed Performance on SciAssess
The detailed results on each tasks in SciAssess are shown in Table 7.
Domain Task SciTulu-7B Mistral-7B Llama3-8B Qwen2-7B SciLitLLM-7B Llama3-70B Qwen2-72B SciLitLLM-72B GPT3.5 GPT4o
Fundamental Science Average 32.3 48.3 58.5 70.3 74.8 70.9 77.1 78.4 62.2 76.7
MMLU 35.5 52.1 59.4 64.0 65.3 76.2 78.7 78.5 64.3 84.2
CMMLU 27.5 31.8 46.1 79.2 87.8 65.5 86.6 89.6 44.9 78.7
Xz-Ch 34.6 51.7 64.7 71.7 75.4 73.4 74.0 75.5 73.2 73.4
Xz-En 31.5 57.6 63.6 66.2 70.7 68.4 69.1 69.9 66.5 70.3
[…]
AlloyQA 6.7 33.3 26.7 53.3 53.3 46.7 53.3 66.7 53.3 46.7
CompEx 9.0 9.0 8.1 10.1 7.2 34.7 19.3 19.5 24.8 50.5
TempEx 34.3 28.5 32.9 28.5 30.9 55.6 58.0 57.9 30.9 60.9
SampDiff 27.4 14.3 29.1 6.3 18.1 26.6 7.6 32.8 12.7 34.6
TreatSeq 42.2 54.9 67.6 65.7 68.6 60.8 75.5 68.6 38.2 67.6
[…]
BioQA 33.3 37.4 43.4 41.4 38.4 50.5 55.6 54.5 29.3 59.6
ChemER 68.3 93.2 84.3 92.1 90.5 86.1 90.9 91.4 93.3 90.3
DisER 80.8 82.2 80.9 87.2 88.4 79.8 81.7 80.9 90.5 81.1
CompDis 67.7 70.3 74.6 73.8 74.5 78.2 76.3 74.0 71.6 72.6
GeneFunc 70.9 79.9 88.8 92.9 89.6 87.1 85.5 81.3 88.8 96.4
GeneReg 85.9 92.8 92.1 97.1 95.9 95.9 95.9 95.7 94.2 93.7
[…]
AffEx 1.1 4.6 5.7 4.3 3.0 3.5 6.1 5.4 8.1 31.4
DrugQA 40.0 53.3 33.3 20.0 46.7 40.0 33.3 40.0 33.3 53.3
TagMol 7.3 0.0 10.2 1.5 20.1 23.0 13.0 27.1 0.6 7.8
MarkMol 28.4 15.9 18.0 34.4 32.8 53.3 31.7 52.4 48.8 63.8
MolDoc 44.0 46.0 50.0 56.0 56.0 48.0 50.0 58.0 44.0 54.0
ReactQA 25.3 25.3 32.6 28.4 26.3 50.5 36.8 32.6 34.7 37.9
ResTarg 31.9 66.2 74.3 77.0 47.7 72.3 77.3 77.3 47.4 55.7
[…]
ElecQA 26.0 20.0 41.0 30.0 28.0 33.0 49.0 33.0 26.0 68.0
OLEDEx 1.8 8.0 6.5 7.2 6.8 16.1 17.4 11.0 13.5 27.9
PolyQA 6.7 6.7 13.3 20.0 93.3 80.0 73.3 80.0 0.0 80.0
PolyCompQA 23.9 25.7 35.8 32.1 49.5 53.2 73.4 74.9 33.0 82.6
PolyPropEx 4.9 20.2 22.4 32.2 43.4 35.9 54.4 54.4 39.5 75.3
SolEx 26.2 31.8 34.0 35.7 32.9 36.2 42.3 38.5 35.8 45.9
ReactMechQA 27.3 31.8 18.2 40.9 18.2 36.4 59.1 48.2 22.7 59.1
Overall Average 33.2 40.6 45.0 48.8 52.4 55.7 57.8 59.5 45.5 63.4
Table 7. Detailed model performance on SciAssess tasks. SciLitLLM-7B shows significant improvement in the Fundamental Science and Organic Materials domains while maintaining comparable performance in other domains. Overall, SciLitLLM-7B achieves approximately 3.6% improvement over the second-best LLM.